OpenStudy (anonymous):

(4A+3B)(-1) Solve. A(x)=|2x-3|. B(x)= x+3 if x<=-1, x^2 if -1

OpenStudy (anonymous):

Wait what are trying to find? @Bladerunner1122

OpenStudy (anonymous):

I assume that A can just be substituted in, but what about B(x) since I don't actually have an x given to me. I'm told to simplify and evaluate.

OpenStudy (anonymous):

@Bladerunner1122 The thing is that I don't know what you're trying to solve for or find.

OpenStudy (anonymous):

Like what's the (4A + 3B)(-1) supposed to denote?

OpenStudy (anonymous):

@Bladerunner1122 hello?

OpenStudy (anonymous):

A and B are the functions

OpenStudy (anonymous):

It says using the definitions of the fuctions, simplify or evaluate (4A+3B)(-1).

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@mathessentials

OpenStudy (jdoe0001):

\(\bf \textit{let us first, do}\quad (4A+3B)\\ A(x) = |2x-3|\qquad \qquad B(x)= x+3\\ 4A\implies 4|2x-3|\qquad \qquad 3B\implies 3(x+3)\\ \quad \\ (4A+3B)\implies 4|2x-3|\quad +\quad 3(x+3)\\ \quad \\ (4A+3B)(-1) = 4|2(-1)-3|\quad +\quad 3((-1)+3)\implies -28+6= -22\\ \quad \\ \textit{so now we can say that}\\ \quad \\ f(x)\implies \begin{cases} -22 & x\le 1\\ \quad \\ \bf x^2& -1<x<2\\ \quad \\ \bf 5 & x=2 \end{cases}\)

OpenStudy (jdoe0001):

anyhow, got a bit truncated... but surely -28+6= -22 :)